OpenStudy (anonymous):

Evaluate the function f(x) = 5 ● 3x for x = -3.

OpenStudy (anonymous):

5/27 ?

OpenStudy (cggurumanjunath):

F(X) = 5 * 3 * X =5 * 3 * - 3 =15 * - 3 = - 45

OpenStudy (anonymous):

my answer is wrong?

OpenStudy (shaadykhamlichi):

yes because x=-3 so f(x)=5.3(-3) f(x)=5.(-9)=-45